OpenStudy (anonymous):

m^2 + 2m - 3/m^2 +4m -5 x m^2 - 3m - 10/m^2 + 5m + 6 i got -3m/10m

OpenStudy (anonymous):

m^2 + 2m - 3/m^2 +4m -5 times m^2 - 3m - 10/m^2 + 5m + 6 i got -3m/10m

OpenStudy (anonymous):

no i don't think that is correct

OpenStudy (anonymous):

neither do i :p

OpenStudy (anonymous):

after factoring numerator is \[(m-1)(m+3)(m-5)(m+2)\]

OpenStudy (anonymous):

denominator is \[(m+5)(m-1)(m+3)(m+2)\]

OpenStudy (anonymous):

oh i didnt factor :|

OpenStudy (anonymous):

gotta factor and cancel.

OpenStudy (anonymous):

common factor are m-1, m+3, m+2

OpenStudy (anonymous):

can you help me with one more?

OpenStudy (anonymous):

10-5x/3 divided by x-2/12x short one

OpenStudy (anonymous):

leaving \[\frac{m-5}{m+5}\]

OpenStudy (anonymous):

\[\frac{10-5x}{3}\div\frac{x-2}{12x}\]?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

just like with regular fraction invert and multiply

OpenStudy (anonymous):

i got 120x-60x^2/3x-6

OpenStudy (anonymous):

\[\frac{10-5x}{3}\times \frac{12x}{x-2}\]

OpenStudy (anonymous):

cancel first, multiply last

OpenStudy (anonymous):

what do i cancel out?

OpenStudy (anonymous):

\[\frac{5(x-2)\times 12x}{3(x-2)}\]

OpenStudy (anonymous):

cancel the 3 and the x - 2

OpenStudy (anonymous):

leaving you with \[5\times 4x=20x\]

OpenStudy (anonymous):

oopsssssssssss

OpenStudy (anonymous):

wrong sorry

OpenStudy (anonymous):

sorry give me a second i factored incorrectly

OpenStudy (anonymous):

cancel the x-2 not the 3

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

\[\frac{10-5x}{3}\times \frac{12x}{x-2}=\frac{5(2-x)\times 12x}{3(x-2)}\]

OpenStudy (anonymous):

i factored incorrectly at first. sorry.

OpenStudy (anonymous):

now 3 goes in to 12 4 times so you can cancel the 3 and replace the 12 by a 4

OpenStudy (anonymous):

yea i saw that but i didnt know cus ur the one with the ansers :p

OpenStudy (anonymous):

and finally \[\frac{2-x}{x-2}=-1\]

OpenStudy (anonymous):

hope that is clear

OpenStudy (anonymous):

so you can cancel the 2-x and the x-2 but make sure to change the sign, because it is -1 and not 1

OpenStudy (anonymous):

whered the 5 go?

OpenStudy (anonymous):

still up there. so you get \[-20x\]

OpenStudy (anonymous):

ohk thanks

OpenStudy (anonymous):

\[5\times 5x\times -1=-20x\]

OpenStudy (anonymous):

typo

OpenStudy (anonymous):

\[5\times 4x\times -1=-20x\]

OpenStudy (anonymous):

so its -20x/x-2?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

x-2 in denominator canceled with 2-x in numerator to give -1

OpenStudy (anonymous):

so there is no denominator at all.

OpenStudy (anonymous):

just -20x

OpenStudy (anonymous):

oh.

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

is it clear that \[\frac{2-x}{x-2}=-1\]?

OpenStudy (anonymous):

i dont see why it turned into -1

OpenStudy (anonymous):

the top is the negative of the bottom

OpenStudy (anonymous):

try it with numbers:\[\frac{2-5}{5-2}=-1\]

OpenStudy (anonymous):

ah. i get it now

OpenStudy (anonymous):

\[\frac{2-10}{10-2}=-1\]

OpenStudy (anonymous):

-3/3 = -1

OpenStudy (anonymous):

-8/8=-1

OpenStudy (anonymous):

a good trick to know\[\frac{a-b}{b-a}=-1\]

OpenStudy (anonymous):

ok well im gettin off now thank you

OpenStudy (anonymous):

yes the top is the negative of the bottom that is why you can cancel

OpenStudy (anonymous):

welcome